OpenStudy (anonymous):

Jordan is a manager of a car dealership. He has two professional car washers, Matthew and Arianna, to clean the entire lot of cars. Matthew can wash all the cars in 14 hours. Arianna can wash all the cars in 11 hours. Jordan wants to know how long it will take them to wash all the cars in the lot if they work together. Write an equation and solve for the time it will take Matthew and Arianna to wash all the cars together. Explain each step.

OpenStudy (anonymous):

This might help http://openstudy.com/study#/updates/55b95fc1e4b0adef802b9a16

OpenStudy (anonymous):

\[\frac{1}{14}+\frac{1}{11}=\frac{1}{x} \]

OpenStudy (anonymous):

1 can be interpreted as the job, the project, the whole thing or all of the cars.

OpenStudy (phi):

one way to think about this problem is use the idea of rate* time = amount of work (a bit like rate*time = distance for figuring how far you go at a certain speed) in these problems, we call the "amount of work" 1 job i.e. 1 without the label. rate is the ratio (or fraction) jobs per hour for example, Matthew can wash all the cars in 14 hours. means Matt does the job in 14 hours or (as a fraction) 1 job per 14 hours or 1 job/14 hours

OpenStudy (phi):

Say we wanted to know how long it takes Matt to do 1 job. Using rate * time = 1 job \[ \frac{1 \ job}{14 \ hours} \cdot t = 1 \ job\] multiply both sides by 14 hours/1 job and simplify to get \[ t = 14 \ hours\] (Which we knew... but this shows how to do the math)

OpenStudy (phi):

with two people, each with their own rate Matt's rate 1 job / 14 hours Arianna's rate 1 job/ 11 hours then we can add their rates, and multiply by the time ("t" because we don't know the number) \[ \left(\frac{1\ job}{14\ hours}+ \frac{1\ job}{11\ hours}\right)t= 1\ job\]

OpenStudy (phi):

You can add the two fractions by using a common denominator of 11*14= 154 can you add \[ \frac{1}{14}+ \frac{1}{11} \] ?

OpenStudy (anonymous):

2/25

OpenStudy (anonymous):

So it's take them 5 hours?

OpenStudy (phi):

oh, I think I see what you did. It's more complicated

OpenStudy (phi):

you need the bottoms of both fractions to be the same number (154 in this case) to get 154 in the first fraction we multiply the bottom by 11, and to keep things from changing we also multiply its top by 11, like this: \[ \frac{1}{14} \cdot \frac{11}{11 } \] (the idea is 11/11 is 1 so it does not change the value, just how it "looks") anyway, multiply top times top, and bottom times top. can you do that ?

OpenStudy (phi):

***anyway, multiply top times top, and bottom times bottom can you do that ?

OpenStudy (anonymous):

so 1 times 14 and 11 times 11?

OpenStudy (phi):

1 * 11 for the tops and 14*11 for the bottom

OpenStudy (phi):

top times top bottom times bottom

OpenStudy (anonymous):

Oh ops. so 11/154

OpenStudy (phi):

ok, now we do the 2nd fraction 1/11 we want its bottom to be 154 what do we multiply top and bottom by ? any idea ?

OpenStudy (anonymous):

multiple it by 14

OpenStudy (phi):

yes, we multiply the 2nd fraction by 14/14 what do we get ?

OpenStudy (anonymous):

14/154?

OpenStudy (phi):

\[ \frac{1}{11} \cdot \frac{14}{14}= \frac{14}{154}\]

OpenStudy (phi):

so we now know \[ \frac{1}{14}+\frac{1}{11} \] is the same as adding \[ \frac{11}{154}+\frac{14}{154} \] when adding fractions with the same bottom, we add their tops and keep the same bottom

OpenStudy (phi):

you should get \[ \frac{11}{154}+\frac{14}{154}= \frac{11+14}{154} = \frac{25}{154} \]

OpenStudy (anonymous):

Yes, I got that but now I don't know what to do after that.

OpenStudy (anonymous):

Do I divide?

OpenStudy (phi):

and your problem is \[ \left( \frac{1}{14}+\frac{1}{11} \right)t = 1 \\ \frac{25}{154}t = 1 \]

OpenStudy (phi):

now multiply both sides by 154/25 and simplify

OpenStudy (anonymous):

so I multiply 25/154 by 154/25?

OpenStudy (phi):

yes, on the left side. and to keep things equal, you multiply the right side by 154/25 like this \[ \frac{154}{25} \cdot \frac{25}{154} t = 1 \cdot \frac{154}{25}\]

OpenStudy (phi):

as you know, when multiplying fractions, you multiply top times top and bottom times bottom on the left side of the equation we do \[ \frac{154 \cdot 25}{25 \cdot 154} \] when multiplying you can change the order. for example, change the order up top: \[ \frac{25 \cdot 154}{25 \cdot 154} \] and we can also write that as \[ \frac{25}{25}\cdot \frac{154}{154} \] notice if you mutiply top times top and bottom times bottom we get back what we started with... this shows it is the same thing. but anything divided by itself is 1 25/25 is 1 and 154/154 is 1 and 1*1 is 1 that is the long way to show that the left side turns into 1*t or just t \[ t = \frac{154}{25} \] you might want to write the right side as a decimal t= 6.16 hours

OpenStudy (anonymous):

Thank you so much!! You were really helpful :)